OpenStudy (anonymous):

could someone please help me? will fan. Which set of side lengths defines a right triangle? A. 7, 9, 13 B. 5, 12, 14 C. 8, 15, 17 D. 9, 12, 14

hartnn (hartnn):

the one which satisfies the Pythagoras theorem

hartnn (hartnn):

\(\Large a^2+b^2 = c^2\) where c = hypotenuse is the largest of the sides

hartnn (hartnn):

so for example, lets take 5,12,14 14 is largest so you will check whether \(14^2\) equals \(5^2+12^2\) or not? \(25+ 144 = 169 \ne 196\) so 5,12,14 does NOT define a right triangle

hartnn (hartnn):

try it for other options

OpenStudy (anonymous):

im working them out on paper so far i don't think it's a either

hartnn (hartnn):

right its not a try c

OpenStudy (anonymous):

i think it's c actually because 8*8=64 and 15*15=225 and when you add those together you get 289 and 17*17 is 289

hartnn (hartnn):

\(\huge \checkmark\) good! :)

OpenStudy (anonymous):

thank you so much for your help!

hartnn (hartnn):

welcome ^_^